OpenStudy (aravindg):

URGENT!!!!!!!!!!11111111

OpenStudy (aravindg):

fIND VECTOR PRODUCT AND MODULUS A=3i+4j B=-3i+7j

OpenStudy (anonymous):

vector product is the determinant of [i j k] [3 4 0] [-3 7 0] = k det (3 4) (-3 7) = k (21+12) = 33k Modulus of 33k is 33.